Question 5. (15 pts) Find the maximum and minimum of S(,y) 22 + y2 = 1. on the circle
Answer #1

Question - 2 f(x,y) xyz The points lie on x² + y2-2 => y² =1-x². f(x, y) = xey² *(1-x2) => f(x, y) x– x3 For maximum and mini2 when n = 1-x² - wit 1- ſ = Ķ 2 2 x = -5, y = 1- * 1-I 23 3 3 Maximum value of flary) xy²=(1/3) - 6) 2 23 - = 313 9 minimum